OpenStudy (anonymous):

The work of a student to find the dimensions of a rectangle of area 6 + 8x and width 2 is shown below: Step 1: 6 + 8x Step 2: 2(4) + 2(6x) Step 3: 2(4 + 6x) Step 4: Dimensions of the rectangle are 2 and 4 + 6x In which step did the student first make an error and what is the correct step? Step 3; 2 + (4 + 6x) Step 3; 2 + (4 ⋅ 6x) Step 2; 2(4) + 2(4x) Step 2; 2(3) + 2(4x)

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

let me read and catch up on that for a bit.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@DeltaForce

OpenStudy (anonymous):

@DarkMoonZ

OpenStudy (anonymous):

@Donblue

OpenStudy (anonymous):

@unicwaan

OpenStudy (anonymous):

@shinebrightlikeadimon

OpenStudy (usukidoll):

whoa... please do not mass tag. It's a form of spam on OS and anyway A = L X W area of the rectangle = length x width

OpenStudy (anonymous):

oh sry

OpenStudy (anonymous):

do u understand the question????

OpenStudy (usukidoll):

so given area 6+8x width 2 6+8x =2L

OpenStudy (usukidoll):

hmmm... perimeter is P=2L+2W but it's not asking that..

OpenStudy (anonymous):

it wants to know where she went wrong

OpenStudy (usukidoll):

area -> 6+8x width -> 2 length -> ? area = width x length 6+8x=2L and supposedly to get length by itself we have to divide 2 3+4x=L

OpenStudy (usukidoll):

so 6+8x = 2(3+4x) ok the width is 2 and the length is 3+4x. student messed up really early.

OpenStudy (anonymous):

ya

OpenStudy (usukidoll):

looks like she is going backwards step 1 we have 6+8x then split these two numbers up 2(3)+2(4x) the 2 is yanked out 2(3+4x) yay! 6x+8 = 2(3+4x) 6x+8=6+8x

OpenStudy (usukidoll):

Step 2; 2(3) + 2(4x) <- this is the correct one. It took a while because I haven't done this in years, but with patience we got what we need ^^

OpenStudy (anonymous):

yup

OpenStudy (usukidoll):

an that part works because I was using A =WL 6+8x=2L and solved for L which was 3+4x=L so by looking at her work carefully I see that she did it differently by not using the formula A =L X W so she split 6 and 8 and factored out a 2 but in her original step 2 she wrote the wrong number

OpenStudy (usukidoll):

sorry stuck in reply mode.

OpenStudy (anonymous):

so step 2 was the wrong one

OpenStudy (usukidoll):

oh yeah we can prove that with ease given that A = 6x+8 and width 2 on her step 2 she had Step 2: 2(4) + 2(6x) which doesn't make sense that 8+12x which is the wrong area

OpenStudy (anonymous):

Step 2; 2(4) + 2(4x) Step 2; 2(3) + 2(4x) which one is right

OpenStudy (usukidoll):

._. the second one

OpenStudy (usukidoll):

distribute the 2 on 2(4)+2(4x) we got the wrong number with x again 8+8x we were NOT given that

OpenStudy (usukidoll):

2(3) + 2(4x) that's 6+8x which is the area we were given in the first place

OpenStudy (usukidoll):

sorry stuck in reply mode.. don't know what's going on ^^

OpenStudy (usukidoll):

but recommendation to student use the formula A = L X W it's easier

OpenStudy (anonymous):

thx soo much

OpenStudy (anonymous):

i will ask more questions later will u be here

OpenStudy (anonymous):

in like 10 min

OpenStudy (usukidoll):

depends. It's the middle of the night where I live... yeah sure. It also depends on the question too...

OpenStudy (anonymous):

easier than this one

OpenStudy (anonymous):

where do u live???

OpenStudy (anonymous):

what country

OpenStudy (usukidoll):

Personal information shouldn't be discussed on OpenStudy... Code of Conduct.

OpenStudy (anonymous):

ohh sorry didnt know

OpenStudy (anonymous):

sorry, i was late to reply to this @3.14192653589